### v1

# Exploring Large-Volume GAGG Scintillators for use in MeV Gamma-Ray Astrophysics

Abstract
Gamma-ray astrophysics in the MeV band is an exciting field in astronomy due to its potential for multi-messenger astrophysics. It has, however, remained under-explored when compared to other wavelengths. One reason for this observational gap is the difficulties with measuring these high-energy photons and the requirement of large amounts of detection material. In this work, we investigate the usage of large-volume GAGG scintillators for use as a calorimeter in future MeV telescopes. We developed a $5\times 5$ array calorimeter utilizing $1\times 1\times 6\ \mathrm{cm}^{3}$ GAGG crystals with C-series SiPM readout. The calorimeter was tested at the High Intensity Gamma-ray Facility (HIGS) with monoenergetic beams ranging from $2-25\ \mathrm{MeV}$. Finally, we also investigated larger $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals and characterized their response across their depth when their surface treatment is either polished or frosted.
keywords:
GAGG, gamma-ray spectroscopy, gamma-ray telescopes, calorimetery, gamma-ray astrophysics
journal: Journal for Radiation Physics and Chemistry
1 Introduction
[…]
Distribution Statement A. Approved for public release: distribution is unlimited.
The gamma-ray astrophysical sky in the MeV range remains largely under-explored. The most recent large-scale mission dedicated to measuring and imaging the gamma-ray sky was the Compton Gamma-Ray Observatory (CGRO), which was de-orbited in 2000 [1]. Since then, the community has proposed several Compton and/or pair-production instruments to address what is known as the ‘MeV gap’. Most pair-production instruments are based on a two-subsystem design: a tracker to trace out the movement of the cascade of particles from the interaction of the gamma ray and a calorimeter to capture the ensuing electromagnetic showers. This tracker-calorimeter design is utilized by current missions such as the Fermi Space Telescope-LAT [2] and AGILE [3]. Missions that are in the concept and development phase include AMEGO-X [4], ASTROGAM [5], SMILE-2+ [6], VLAST [7], and GECCO [8].
This work investigates the use of gadolinium aluminum gallium garnet (GAGG) based scintillator to act as a high-energy calorimeter for potential usage in gamma-ray telescopes. GAGG presents several advantages over CsI, as outlined in Tab. 1. First, it has superior energy resolution. Next, it is significantly denser than most other scintillators, which results in a more compact design. A faster decay time also allows for operation at higher count rates.
Although GAGG is a relatively new material, it does have a minor space heritage. Launched in December 2021, the GAgg Radiation Instrument (GARI) 1A and 1B [9] have been operating for several years [10]. Next, there have been several space-based instruments proposed that utilize GAGG: GALI [11], HERMES [12],LISSAN [13], GRAPE [14], and FGS [15]. In addition, several studies have measured the internal activation and degradation associated with high-energy proton exposure [16, 17, 18].
Table 1: Some scintillator characteristics comparing GAGG and CsI:Tl. The bold text signifies the better crystal for the given property.
GAGG CsI:Tl
Light Yield (ph/keV) 45 54
Resolution (FWHM @ 662 keV) $\sim$5$\sim$8%
Density (g/$\mathrm{cm}^{3}$) 6.3 4.51
Decay Time (ns) $\sim$140 1000
The outline of this manuscript is as follows: Sec. 2 describes the development and construction of a $5\times 5$ array GAGG calorimeter utilizing $1\times 1\times 6\ \mathrm{cm}^{3}$ crystals. Sec. 3 presents the calorimeter’s response to high-energy gamma rays ranging from 2 MeV to 25 MeV. Finally, Sec. 4 investigates longer crystals $(1\times 1\times 8\ \mathrm{cm}^{3})$ with different surface treatments and their response as a function of depth. A presents the simulated response of the $5\times 5$ array calorimeter to high-energy gamma rays.
2 Development of a $5\times 5$ array GAGG Calorimeter
In this work, we utilize high light-yield GAGG. The prototype in this section uses $1\times 1\times 6\ \mathrm{cm}^{3}$ crystals, each one wrap with a diffuse white reflector (Tetratex) [19], and place a $6\times 6\ \mathrm{mm}^{2}$ C-series SiPM by onsemi [20] on each end of the crystal. Fig. 1 shows the bare crystal and a fully developed crystal element.
Figure 1: (Top) shows the bare unwrapped GAGG crystal while (Bottom) shows the fully wrapped and assembled crystal detector element. This crystal in this picture has a volume of $1\times 1\times 6\ \mathrm{cm}^{3}$.
Since the crystals are not long enough to develop a hodoscopic design (as done with the Fermi-LAT [2]), we instead opt for a ‘finger’ based design and place the crystals upright and length-wise along the axis of the telescope. The crystals are housed in a 3D printed fixture arranged in a $5\times 5$ array such that the center-to-center pitch is $1.1\ \mathrm{cm}$. Fig. 2(a) shows the computer-aided design (CAD) model of the calorimeter, while Fig. 2(b) shows the top view of the calorimeter. All SiPMs were readout by an IDE AS ROSSPAD [21]. The entire system is then placed in a box to maintain light tightness, shown in Fig. 3. The general implementation, front-end, and analysis techniques are duplicated from a past prototype of the ComPair CsI subsystem [22].
(a) CAD Model
(b) Top view of the calorimeter
Figure 2: Images of the $5\times 5$ array GAGG calorimeter
Figure 3: The $5\times 5$ array calorimeter is placed inside the test box. The calorimeter array is housed in the upper right (black, 3D printed structure), while the ROSSPAD readout is placed on the left. A custom power supply, placed on the lower left, is utilized as a legacy from a different system that requires a higher current limit.
2.1 Gamma-ray response
The deposited energy is calculated as the geometric mean between each SiPM on each end for a given crystal:
$E=\sqrt{Q_{top}\times Q_{bottom}},$ (1)
where $Q_{top}$ and $Q_{bottom}$ are the measured signals from the top and bottom SiPM on the same crystal.
With the fully assembled calorimeter, we tested its response to ${}^{22}\mathrm{Na},^{137}\mathrm{Cs}$ and Thoriated welding rods. Fig. 4 plots the percent full-width at half-maximum (FWHM) energy resolution as a function of the energy deposited. In the plot, only single crystal spectra are considered (i.e. coincidence between multiple crystals is ignored). The error bars represent the standard deviation between all the bars.
The measured resolution at 662 keV is around 8%, which is poorer than the general advertised value. This is most likely due to the difference in geometry as this GAGG is longer than mainstream crystals that are on the market.
Figure 4: The average FWHM resolution for a given energy for all crystals with the error bars representing the observed standard deviation between the crystals. The sources we used included ${}^{22}\mathrm{Na},^{137}\mathrm{Cs}$ and Thoriated welding rods.
3 High-Energy Gamma-Ray Response of the Calorimeter
In April 2022, the calorimeter was taken to the Triangle Universities Nuclear Laboratory (TUNL) High Intensity Gamma-Ray Source (HIGS) facility. The facility can produce a monoenergetic gamma-ray beam with a diameter of $0.5\ \mathrm{cm}$. The experiment utilized beam energies of 2, 5, 7, 15, and 25 MeV. Fig. 5 shows the experimental setup. Since this experiment was conducted as an add-on to the AMEGO/ComPair campaign [23], the GAGG is placed downstream of the ComPair instrument. The ComPair instrument consists of 3 subsystems: 6 layers of $0.5\ \mathrm{mm}$ of Si, $2\ \mathrm{cm}$ of CZT, and $8.35\ \mathrm{cm}$ of CsI.
Figure 5: Experimental setup of the GAGG array conducted at the HIGS facility. The ComPair instrument is present and in front of the GAGG as this effort was conducted as a piggy back to their campaign.
During the experiment, we placed the calorimeter in different positions and angles to the beam. This is to change the interaction profile of the gammas with the calorimeter. Fig. 6 shows the summed response to different irradiation configurations of a 5 MeV beam. Side irradiation refers to when the beam is shot into the center column of the calorimeter while front irradiation is where the beam is aimed into the $1\times 1\ \mathrm{cm^{2}}$ face of the center crystal in the array. We note that the difference in the full calorimeter response is relatively invariant in that the ratio between the escape peaks, continuum, and full-energy peak are very similar. This is likely due to the near-cubic geometry of the calorimeter which results in a uniform sensitivity to source location.
Figure 6: Full calorimeter response to a 5 MeV beam while the calorimeter is oriented at different poses.
Fig. 7 shows the crystal-by-crystal response to different irradiation configurations relative to the 5 MeV beam. Fig. 7(a) plots the front irradiation while Fig. 7(b) plots side irradiation. The two irradiations show the expected full energy and escape peaks in the targeted crystals, with the surrounding crystals showing significant continuum and 511 keV depositions from the ensuing cascades. Also detected is the 1.46 $\mathrm{MeV}$ from 40K present in the background. When comparing to the summed spectra in Fig. 6, we observe a very different peak-to-escape ratio with the double escape peak being the most prominent feature.
(a) Face irradiation
(b) Side irradiation
Figure 7: Crystal-to-crystal response to 5 MeV beam while the calorimeter is oriented at different poses.
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peak, likely due to the low full-energy efficiency and full-containment of the ensuing spallation.
Figure 8: GAGG Calorimeter response to monoenergetic gamma rays ranging from $2\ \mathrm{MeV}-25\ \mathrm{MeV}$
4 Depth of Interaction response study of GAGG Crystals
We expand our investigation of GAGG by exploring longer crystals that are $1\times 1\times 8\ \mathrm{cm}^{3}$ in volume. Moving to the longer crystals is of interest as it increases the detection efficiency and full-energy containment. In Fermi’s LAT calorimeter, all the sides were roughened to introduce a depth-dependent light output when read out on each end of the crystal. In this section, we investigate the difference in depth of interaction response when the sides of the crystals are either polished or frosted. Fig. 9 shows the roughened crystal (on the right), next to a polished version (left). Each crystal is packaged and readout like the crystals described in Sec. 2.
Figure 9: Picture of the $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals that are either polished (left) or frosted on all sides (right).
4.1 Collimated Depth Measurements
In this section, we conducted collimated measurements at each depth to verify the scintillation properties across the entire crystal and the depth reconstruction. A fan beam is produced by using two lead bricks separated by an American penny with a ${}^{137}\mathrm{Cs}$ source placed on the other end. The assembly is mounted on a robotic gantry that is programmed to move in $5\ \mathrm{mm}$ steps and scan along the length of the GAGG crystal. Fig. 10 illustrates the experimental setup.
Figure 10: Diagram of the experimental setup to measure the depth response of the $1\times 1\times 8\ \mathrm{cm}^{3}$ GAGG crystal.
Fig. 11 plots the ${}^{137}\mathrm{Cs}$ photopeak response of the SiPM on each end. We observe that at the center of the crystal ($40\ \mathrm{mm}$), the signals are not perfectly equal. This could arise from asymmetric gain between the two ends or the difference between the optical coupling of the SiPM to the crystal. Most of the crystal displays a linear response along the depth. Finer steps are required to be taken to observe the edge effect when the interactions are close to one end of the crystal.
Figure 11: Photopeak response of each SiPM to collimated ${}^{137}\mathrm{Cs}$ source placed at different depth along the crystal.
Like in the Fermi LAT Calorimeter, we calculate the depth of interaction (DOI) using Eq. 2.
$\mathrm{DOI}=\frac{Q_{top}-Q_{bottom}}{Q_{top}+Q_{bottom}}.$ (2)
We can then make a DOI calculation as a function of the collimator position and plot it in Fig. 12. The figure also plots the calculated DOI for a polished crystal. The frosted crystal displays a very linear response to the different depths indicating that applied surface roughening is sufficient. On the other hand, the polished crystal displays a negligible amount of depth response. Nevertheless, both crystals display a fully active scintillator.
Figure 12: Calculated DOI as a function of collimator location for a crystal that was either polished or frosted on each end.
4.2 Flood Irradiation Response
This section demonstrates the response of 3 frosted and 3 polished crystals to a ${}^{137}\mathrm{Cs}$ flood irradiation. Using Eq. 1 and 2, we can develop a bivariate histogram to study the depth-energy response. Fig. 13 plots the depth-energy response of the frosted (top row) and polished crystals (bottom row). Once again, the polished crystals display no depth response while the frosted ones show a depth response. In the frosted crystals, the photopeak is visible along the multiple depths. The DOI value range is not centered around 0 for all crystals, likely due to the asymmetric gain between each end.
Figure 13: Raw depth-energy response to a ${}^{137}\mathrm{Cs}$ source. The top row plots the 3 different frosted crystals while the bottom row plots the 3 responses of the polished crystals.
The energy and depth are then calibrated using techniques described in [22]. Fig. 14 plots the corrected response, highlighting the ${}^{137}\mathrm{Cs}$ peak, where the top and bottom rows plot the frosted and polished crystals respectively.
Figure 14: Corrected Depth Energy response of the $1\times 1\times 8\ \mathrm{cm}^{2}$ GAGG crystals. The top row plots the 3 different frosted crystals while the bottom row plots 3 polished crystals.
5 Conclusion and Future Work
This work explores the usage of GAGG as a high-energy calorimeter for a MeV range gamma-ray telescope. We built a $5\times 5$ array of GAGG crystals, each $1\times 1\times 5\ \mathrm{cm}^{3}$ in size. This system was then characterized with monoenergetic gamma rays ranging from $2$ to $25\ \mathrm{MeV}$. The results show good agreement with simulations. Next, we investigate the depth response of $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals that are either frosted or polished on all sides. Collimated measurements show that the longer $8\ \mathrm{cm}$ crystals are fully active and display a near-linear response when using standard DOI calculations.
We are currently developing a larger $8\times 8$ array calorimeter using the $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals. In addition, we are developing SiPM readouts that field two different SiPM species such that their different gains will the dynamic energy range of the readout [24]. The ‘dual-gain’ SiPMs are currently under fabrication by Fondazione Bruno Kessler [25].
Acknowledgment
This work is supported by the Office of Naval Research 6.1. We are grateful for the efforts of the staff at the HIGS facility to ensure a successful experiment.
We thank the AMEGO/ComPair team for the opportunity to join their HIGS Campaign. ComPair is funded under NASA Astrophysics Research and Analysis (APRA) grants NNH14ZDA001N- APRA, NNH15ZDA001N-APRA, NNH18ZDA001N-APRA, NNH21ZDA001N-APRA.
Appendix A Simulated Response of the GAGG Calorimeter
We modeled the $5\times 5$ GAGG calorimeter using the SWORD simulation package [26]. In addition to the calorimeter (shown in Fig. 2(a)), the ComPair system was also modeled to simulate the HIGS beam test conditions. The plots in this section show the GAGG calorimeter’s response to $2,5,7,15$ and $25\ \mathrm{MeV}$ monoenergetic gamma rays. The simulated energy blurring is accomplished using the mean energy resolution presented in Fig. 4. However, we did not account for the crystal to crystal variations in energy response.
Figure 15: CAD model of the simulated environment and instrument mass model for the experimental setup taken place at HIGS.
For energies ranging from $2-7\ \mathrm{MeV}$, a clear full energy peak is visible along with the two escape peaks, which mirrors what is observed during experimentation. The $15\ \mathrm{MeV}$ simulated run shows very modest full energy and escape peaks. On the other hand, the experimental result shows a skewed left continuum with no obvious features. This is likely a result of the simulation not accounting for the crystal to crystal variations. The $25\ \mathrm{MeV}$ run shows a continuum with no spectral peaks, similar to the experimental run.
Figure 16: Simulated response of the 5x5 array GAGG calorimeter to 2 MeV gamma rays.
Figure 17: Simulated response of the 5x5 array GAGG calorimeter to 5 MeV gamma rays.
Figure 18: Simulated response of the 5x5 array GAGG calorimeter to 7 MeV gamma rays.
Figure 19: Simulated response of the 5x5 array GAGG calorimeter to 15 MeV gamma rays.
Figure 20: Simulated response of the 5x5 array GAGG calorimeter to 25 MeV gamma rays.